25. When EP = -2, the optimal markup on price is:

a.

100%

b.

67%

c.

50%

d.

33%
Answer #1

MR=P*(1+(1/Ep))

Optimal condition is achieved when MR=MC

MC=P*(1+(1/Ep))

Price=MC(1+Markup)

1=(1+Markup)*(1+(1/Ep))

(1+(1/Ep))^-1=(1+Markup)

(1+(1/Ep))^-1-1=Optimal Markup

(1+(1/-2))^-1-1=Optimal Markup

Optimal Markup =1=100%

Option A is correct